Evaluate using root test.

Infinity

Summation sign (1- 1/k)^k

k = 1
1
Expert's answer
2020-02-23T15:11:13-0500

"lim _{k\to \infty}((1-1/(k+1))^{k+1}/(1-1/k)^k)"

"lim _{k\to \infty}((k/(k+1))^{k+1}/((k-1)/k)^k)"

"lim _{k\to \infty}((k^2/(k^2-1))^{k})"

"lim _{k\to \infty}((1/(1-1/k^2))^{k})" is of the form "1^{\infty}"

"e^{lim _{k\to \infty}(((k^2/(k^2-1))-1)k)}"

"e^{lim_{k \to \infty}(k/(k^2-1))}"

"e^0=1"

Root test fails as limit value is equal to 1.